# Present address: Physics Dept., U.C. Santa Barbara, Santa Barbara, California 93106, USA and Google Quantum AI, 301 Mentor Dr, Goleta, CA 93111, USA

### v2

@@ Abstract
Ionizing radiation impacts create bursts of quasiparticle density in superconducting qubits. These bursts temporarily degrade qubit coherence which can be detrimental for quantum error correction. Here, we experimentally resolve quasiparticle bursts in 3D gap-engineered transmon qubits by continuously monitoring qubit transitions. Gap engineering allowed us to reduce the burst detection rate by a factor of a few. This reduction falls several orders of magnitude short of that expected if the quasiparticles were to quickly thermalize to the cryostat temperature. We associate the limited effect of gap engineering with the slow thermalization of the phonons in our chips after the burst.
@@
Introduction.—Quantum error correction protocols redundantly encode a logical qubit in the entangled state of multiple physical qubits [nielsen_quantum_2010], making encoded information robust to errors on individual qubits. This holds if errors are uncorrelated between different physical qubits and lack temporal correlations on the same qubit. Common errors in superconducting processors (e.g. qubit relaxation from material defects) generally satisfy these assumptions. However, there are rare error burst events that violate them. During these error bursts (hereby referred to as “bursts”), qubit relaxation rates temporarily increase across multiple qubits simultaneously [vepsalainen_impact_2020, martinis_saving_2021, wilen_correlated_2021, mcewen_resolving_2022, thorbeck_two-level-system_2023, harrington_synchronous_2024, li_direct_2024]. Mitigating bursts is therefore essential to realize large-scale quantum error correction.
The underlying mechanism of the bursts is the impact of ionizing radiation on the chip hosting the qubits  [^1] There is another mechanism for bursts, where a phonon shower is generated by the avalanche-like release of stress in the chip [yelton_correlated_2025]. Its experimental signatures and mitigation strategies should be similar to those for high-energy impacts.. A high-energy particle striking the chip produces a cascade of high-energy phonons that rapidly propagate throughout the substrate. When the phonons collide with superconducting films of the qubits, they break Cooper pairs, producing quasiparticle excitations (QPs). When the QPs tunnel across the Josephson junction, they can exchange energy with the qubit degree of freedom. This results in an elevated qubit relaxation rate. Only when the QPs recombine or get trapped does the error burst end and the processor performance recover. This prompts the search for strategies to either accelerate the phonon decay [martinis_saving_2021, iaia_phonon_2022] or suppress QP tunneling.
@@
In this paper, we address this question by measuring QP bursts in 3D transmons with different pairs of junction film thicknesses and, consequently, with different gap profiles. First, we benchmark our gap engineering by measuring the gap difference, $\delta\Delta$, for our devices with different film thicknesses. The gap difference is obtained by monitoring the tunneling rate of resident QPs at different fridge temperatures. By varying the film thickness, we control the gap difference from 0 to 10 $\mathrm{GHz}$ (in frequency units). We then evaluate the effect of gap difference on the detection rate of burst events and measure qubit relaxation rates during bursts.
@@
We find that gap engineering reduces the burst detection rate by a factor of a few. However, the reduction falls short of the four orders of magnitude suppression observed for the resident QPs. We believe that this discrepancy results from the elevation of the effective chip temperature during the burst to about 90 mK. The additional thermal energy allows the QPs to overcome the gap difference and tunnel across the junction, thus facilitating qubit relaxation. We directly confirm the persistence of elevated substrate temperature by comparing the qubit excitation and relaxation rates during the burst.
@@
Figure 1: (a) Parity-switching rate ($\Gamma$) as a function of the fridge temperature ($T$) for devices with different film thicknesses. The film thickness is used as a knob to control the gap difference $\>\delta\Delta$ at the qubit junction. Down (up)-pointing triangles represent rate $\Gamma_{0}$ ($\Gamma_{1}$) conditioned on the qubit residing in the ground (excited) state. When $\>\delta\Delta>hf_{q}$, the gap difference is efficient in obstructing the tunneling of resident QPs. Correspondingly, $\Gamma_{1}$ follows an Arrhenius law with the activation exponent $\>\delta\Delta-hf_{q}$, i.e., $\Gamma_{1}\propto\exp\left(-\left[\>\delta\Delta-hf_{q}\right]/\>k_{\mathrm{B}}T\right)$. This regime is realized in the big- and medium-$\>\delta\Delta$ device (left and middle panels). For the small-$\>\delta\Delta$ device (right panel), $\>\delta\Delta<hf_{q}$, and the gap difference is not effective at preventing the tunneling of resident QPs. The observed rate $\Gamma_{1}$ for the small-$\>\delta\Delta$ device is the highest among the measured devices. Comparing the data for $\Gamma_{1}$ and $\Gamma_{0}$ to theory (solid lines) allows us to extract $\>\delta\Delta$ and the resident QP density for all three devices. (b) Exponential suppression of the resident QP tunneling rate, $\Gamma_{1}-\Gamma_{1}^{\mathrm{ph}}$, in state $|1\rangle$. Different points correspond to devices with different $\>\delta\Delta-\>hf_{q}$, different cooldowns, and temperatures [noauthor_see_nodate]. For convenience of comparison between different devices, the rate is normalized by the resident QP density, $\>\gamma^{\mathrm{QP}}_{1}\coloneqq(\Gamma_{1}-\Gamma_{1}^{\mathrm{ph}})/\>x_{\mathrm{QP}}^{\mathrm{ne}}$. When the gap difference surpasses the qubit energy, the normalized tunneling rate decreases exponentially with increasing $\>(\delta\Delta-hf_{q})/k_{\mathrm{B}}T$. The inset shows the full dependence of $\>\gamma^{\mathrm{QP}}_{1}$ on $\>(\delta\Delta-hf_{q})/k_{\mathrm{B}}T$. The normalized rate diverges when $\>\delta\Delta=\>hf_{q}$ but becomes temperature-insensitive for $\>\delta\Delta<\>hf_{q}$. The gray-shaded regions in the main plot and inset represent the theoretical prediction for the rate for a range of additional parameters such as $E_{J}$ or $E_{C}$.
@@
Our results should be contrasted with Ref. [mcewen_resisting_2024], where the gap difference – obtained by film thicknesses similar to our gap-engineered device – resulted in essentially no bursts being detected. The difference may arise from a combination of several factors. First, unlike ours, the chip in Ref. [mcewen_resisting_2024] has most of its area covered by the aluminum ground plane. Second, their substrate is silicon whereas ours is sapphire. Third, the chips are thermalized differently. Our chips are anchored to the package only by two isolated clamps each, whereas Ref. [mcewen_resisting_2024] presumably used numerous wirebonds that anchor their chip. Finally, unlike in our measurement, Ref. [mcewen_resisting_2024] probed the qubit only sparsely, once every 100 $\mathrm{\mu s}$. Such a measurement may miss bursts recovering faster than 100 $\mathrm{\mu s}$.
@@
Experimental setup.— Our goal is to evaluate the effectiveness of gap engineering in suppressing decoherence caused by the QP bursts. The gap difference between the junction leads, with gaps of $\Delta$ and $\Delta+\>\delta\Delta$, prevents the tunneling of low-energy QPs from the low-gap lead to the high-gap one. The elevation of the superconducting gap in one of the leads, $\delta\Delta$, is controlled by changing the film thicknesses employed in the junction fabrication [chubov_dependence_1969, yamamoto_parity_2006, court_energy_2007, diamond_distinguishing_2022, marchegiani_quasiparticles_2022, connolly_coexistence_2024]. We fabricate three 3D aluminum transmons using different film thickness pairs: [17 nm, 83 nm], [23 nm, 123 nm], and [85 nm, 106 nm] (measured by atomic force microscopy). Based on prior measurements [chubov_dependence_1969, yamamoto_parity_2006, court_energy_2007, marchegiani_quasiparticles_2022], the corresponding $\>\delta\Delta$ are expected to be approximately 10 GHz (“big”-$\>\delta\Delta$), 5 GHz (“medium”-$\>\delta\Delta$), and 0.5 GHz (“small”-$\>\delta\Delta$), respectively, in frequency units. Table 1 summarizes relevant qubit parameters.
@@
The transmons are designed to be in the offset-charge-sensitive regime [riste_millisecond_2013, serniak_hot_2018, serniak_direct_2019], with the charge dispersion on the order of MHz. This enables the detection of charge-parity-switching events, which are experimental signatures of QP tunneling. By measuring the parity-switching rates, we can monitor how frequently QPs cross the junction in the presence of $\>\delta\Delta$. All three chips are mounted together in a multiplexed package, ensuring that electromagnetic environment is the same [noauthor_see_nodate].
@@
Table 1: Fit parameters and uncertainties corresponding to devices in Fig. 1(a). For the medium- and small-$\>\delta\Delta$ devices, the parity-switching rate is dominated by the QP tunneling rate, making it difficult to extract $\Gamma_{1}^{\mathrm{ph}}$. Therefore, for these devices, we assume $\Gamma_{0}^{\mathrm{ph}}=\Gamma_{1}^{\mathrm{ph}}$. In contrast, we can confidently fit $\Gamma_{1}^{\mathrm{ph}}$ for the big-$\>\delta\Delta$ device. This value can be found in [noauthor_see_nodate].
$f_{q}$ $\>\delta\Delta/h$ $\Delta/h$ $\>x_{\mathrm{QP}}^{\mathrm{ne}}$ $\Gamma_{0}^{\mathrm{ph}}$
(GHz) (GHz) (GHz) (×$\times 10^{-10}$) ($\mathrm{s}^{-1}$)
big-$\>\delta\Delta$ 3.57 8.2$\pm$0.4 45.2$\pm$0.4 6.3$\pm$1.8 0.22$\pm$0.01
med.-$\>\delta\Delta$ 4.78 5.8$\pm$0.1 46.1$\pm$0.1 1.0$\pm$0.1 0.24$\pm$0.01
small-$\>\delta\Delta$ 4.27 0$\pm$1.1 45.1$\pm$0.4 1.9$\pm$0.1 0.52$\pm$0.05
@@
Measuring the gap difference.—As an initial step, we characterize the superconducting gap barriers, $\delta\Delta$, of our devices. To do this, we measure the parity-switching rates in the steady state, i.e., away from rare QP bursts. In this case, the parity-switching rates are dominated by the tunneling of resident non-equilibrium QPs populating superconducting films. The energy distribution of these QPs equilibrates with the base stage of the refrigerator [connolly_coexistence_2024]. The temperature dependence of the QP tunneling rate then follows the Arrhenius law, with the activation exponent controlled by $\>\delta\Delta$. Indeed, $\>\delta\Delta$ impedes the QPs across the junction, such that only few energetic QPs tunnel through the junction. Fitting the measured temperature dependence of the parity-switching rates to the Arrhenius law thus directly yields $\>\delta\Delta$, as we explain below. This measurement resembles that in Ref. [connolly_coexistence_2024]. Figure 1(a) shows the measured parity-switching rates as a function of temperature for our three devices.
@@
To fit the measured rates, we use a detailed theoretical model of QP tunneling in the presence of a gap difference at the junction [diamond_distinguishing_2022, marchegiani_quasiparticles_2022, connolly_coexistence_2024]. According to this theory, when the qubit is in the ground state, the parity-switching rate reads $\>\Gamma^{\mathrm{QP}}_{0}=\>\gamma^{\mathrm{QP}}_{0}\>x_{\mathrm{QP}}^{\mathrm{ne}}$. Here, $\>x_{\mathrm{QP}}^{\mathrm{ne}}$ is the temperature-independent density of resident QPs (in units of Cooper pair density), and $\>\gamma^{\mathrm{QP}}_{0}\propto E_{J}\exp\left(-\>\delta\Delta/\>k_{\mathrm{B}}T\right)$ is proportional to the tunneling rate and the fraction of QPs with energy exceeding $\delta\Delta$. Here, $\>k_{\mathrm{B}}$ is the Boltzmann constant, $T$ is the QP temperature, and $E_{J}$ is the Josephson energy of the junction. When the qubit is excited, and if $\>\delta\Delta>\>hf_{q}$, this expression modifies to $\>\Gamma^{\mathrm{QP}}_{1}=\>\gamma^{\mathrm{QP}}_{1}\>x_{\mathrm{QP}}^{\mathrm{ne}}$ with $\>\gamma^{\mathrm{QP}}_{1}\propto E_{J}\exp\left(-\left[\>\delta\Delta-\>hf_{q}\right]/\>k_{\mathrm{B}}T\right)$, where $h$ is the Planck constant and $f_{q}$ is the qubit frequency. The activation energy is reduced compared to that in $\>\gamma^{\mathrm{QP}}_{0}$ because the QP can absorb the excitation from the qubit, making it easier to overcome the gap difference. On contrary, if $\>\delta\Delta<\>hf_{q}$, then $\>\gamma^{\mathrm{QP}}_{1}$ becomes roughly temperature-independent. For each device, we simultaneously fit $\Gamma_{0}$ and $\Gamma_{1}$ to the expressions outlined above, using $\>\delta\Delta$ and $\>x_{\mathrm{QP}}^{\mathrm{ne}}$ as fitting parameters.
@@
There are two complications in this fitting protocol. First, at high temperatures, $T>100\>\mathrm{mK}$, the density of equilibrium QPs, $\>x_{\mathrm{QP}}^{\mathrm{th}}(T)=\sqrt{2\pi\>k_{\mathrm{B}}T/\Delta}\exp\left(-\Delta/\>k_{\mathrm{B}}T\right)$, becomes comparable to that of the resident non-equilibrium QPs, $\>x_{\mathrm{QP}}^{\mathrm{ne}}$. Therefore, at high temperatures, we account for the contribution of equilibrium QPs to the parity-switching rates. Second, there is a small background contribution $\Gamma^{\mathrm{ph}}$ to parity-switching rates coming from the absorption of stray pair-breaking photons at the junction [houzet_photon-assisted_2019, diamond_distinguishing_2022, connolly_coexistence_2024]. In our theoretical fit, we add such contributions as a temperature-independent offset. Notably, the contribution $\Gamma^{\mathrm{ph}}$ of the photons might differ for the two qubit states [houzet_photon-assisted_2019, diamond_distinguishing_2022]. We use notations $\Gamma_{0}^{\mathrm{ph}}$ and $\Gamma_{1}^{\mathrm{ph}}$ to differentiate this contribution between the respective qubit states.
Figure 2: (a) Schematic of how a high-energy impact leads to a QP burst. The collision of high-energy particles ionizes electron-hole pairs in the substrate. When the electron-hole pairs recombine, they generate high-energy phonons (green waves). These phonons have enough energy to break Cooper pairs in superconducting films, thereby generating numerous QPs. Tunneling of these QPs across the junction increases the qubit relaxation rate. (b) Pulse sequence for detecting QP bursts. Every 5.7 $\mathrm{\mu s}$, the qubit is read-out, then projected into |$\ket{1}$. (c) Histogram of the qubit relaxation event counts within a 1 ms window. Different colors correspond to different devices. In the absence of bursts, we expect the histograms to represent a Poisson process. The most likely number of relaxation events is determined by the steady-state coherence time $T_{1}^{\rm steady}$. The positions of the observed maxima are consistent with measured steady-state relaxation times ($T_{1}^{\rm steady}\approx 100\mathrm{\mu s}$). We associate clear deviation from the Poissonian statistics in the tails of the distributions with the rare QP bursts. To the right of the black line, the measured counts for both devices are dominated by QP bursts. Post-selecting events based on this threshold thus allows fair comparison of the burst occurrence rate in the big-$\>\delta\Delta$ and small-$\>\delta\Delta$ devices. The larger $\>\delta\Delta$ yields roughly five times fewer detected burst events according to this criteria. (d) A readout trace during a QP burst. After the burst begins, the qubit experiences an excess number of relaxation events. During the burst, the number of relaxation events in several 1 ms bins exceeds the threshold in panel (c). It allows us to identify the burst. (e) Excess qubit relaxation rate $\Delta\Gamma_{10}$ after high-energy impacts, averaged over all observed burst events. The bursts are identified according to the threshold defined in panel (c). Exponential fit results are shown with solid lines. The extracted burst duration, see the legend, is insensitive to $\>\delta\Delta$. The inset shows the qubit temperature, $T_{q}$, during bursts in the big-$\>\delta\Delta$ device (see [noauthor_see_nodate] for the measurement protocol). We observe $T_{q}$ abruptly increases from the background value (50 mK) to $\approx 90$ mK and stays consistently elevated for over 5 ms. We attribute this increase to elevated QP temperature during the burst. We note that the 50 mK background likely does not reflect the real temperature of the sample. Excess of |$\ket{1}$ measurements primarily stems from measurement errors.
@@
Burst measurement.—With the gap differences under control, we proceed to analyze the temporally correlated errors caused by QP bursts in the devices. To assess the effect of the gap difference on mitigating the high-energy impacts, we compare the bursts in the big- and small-$\>\delta\Delta$ devices. The results for the medium-$\>\delta\Delta$ device are discussed separately in [noauthor_see_nodate], due to the complications arising from its gap difference being comparable to the qubit energy.
@@
To detect the QP bursts, we attempt to stabilize the qubit in its excited state. To this end, we repeatedly readout the qubit state every 5.7 $\mathrm{\mu s}$. If we find the qubit in its ground state $|0\rangle$, we apply a $\pi$-pulse to bring it to the excited state $|1\rangle$. The pulse sequence is illustrated in Fig. 2(b). During QP burst events, the qubit relaxation rate significantly increases. This allows us to identify the bursts by the excess number of observed $|1\rangle\rightarrow|0\rangle$ relaxation events [e.g. Fig. 2(d)]. We note that the measurement is post-selected for events not causing large changes to offset charge [noauthor_see_nodate]. However, we find that this post-selection is inconsequential for the burst event statistics.
To collect sufficient statistics, we repeat the measurement for about 10 hours for each device. We then chop the measurement traces into $1\>\mathrm{ms}$ windows and tally the number of observed relaxation events in each window. Normalized histograms of the relaxation counts are shown in Fig. 2(c). For each device, the histogram contains a large peak (10-15 relaxation events). This peak is unrelated to the QP bursts. It describes a Poissonian relaxation process consistent with the independently measured steady-state qubit relaxation time $T_{1}^{\rm steady}\sim 100\>\mu\mathrm{s}$. In addition to the Poissonian peak, however, there is an anomalous tail with an excess number of relaxation events, which stems from the bursts. Its clear distinction from the Poissonian peak makes us confident that a burst has occurred within a $1\>\mathrm{ms}$ window when the number of relaxation events exceeded a certain threshold  [^2] To increase the confidence in identifying the bursts, we additionally rely on temporal correlation of the relaxation events and reject sequences where the qubit has leaked to more highly excited states, see [noauthor_see_nodate] for details. Choosing a common threshold for both devices, we see that the burst detection rate in the big-$\delta\Delta$ device ($0.57\>\mathrm{min}^{-1}$) is lower than in the small-$\>\delta\Delta$ device ($2.76\>\mathrm{min}^{-1}$).

### v3

@@ Abstract
Ionizing radiation impacts create bursts of quasiparticle density in superconducting qubits. These bursts temporarily degrade qubit coherence which can be detrimental for quantum error correction. Here, we experimentally resolve quasiparticle bursts in 3D gap-engineered transmon qubits by continuously monitoring qubit transitions. Gap engineering allows us to reduce the burst detection rate by a factor of five. This reduction falls four orders of magnitude short of that expected if the quasiparticles were to quickly thermalize to the cryostat temperature. We associate the limited effect of gap engineering with the slow thermalization of the phonons in our chips after the burst.
@@
Introduction.—Quantum error correction protocols redundantly encode a logical qubit in the entangled state of multiple physical qubits [nielsen_quantum_2010], making encoded information robust to errors on individual qubits. This holds if errors are uncorrelated between different physical qubits and lack temporal correlations on the same qubit. Common errors in superconducting processors (e.g. qubit relaxation from material defects) generally satisfy these assumptions. However, there are rare error burst events that violate them. During these error bursts (hereby referred to as “bursts”), qubit relaxation rates temporarily increase across multiple qubits simultaneously [vepsalainen_impact_2020, martinis_saving_2021, wilen_correlated_2021, mcewen_resolving_2022, thorbeck_two-level-system_2023, harrington_synchronous_2025, li_cosmic-ray-induced_2025]. Mitigating bursts is therefore essential to realize large-scale quantum error correction.
The underlying mechanism of the bursts is the impact of ionizing radiation on the chip hosting the qubits  [^1] There is another mechanism for bursts, where a phonon shower is generated by the avalanche-like release of stress in the chip [yelton_correlated_2025]. Its experimental signatures and mitigation strategies should be similar to those for high-energy impacts.. As illustrated in Fig. 2(a), a high-energy particle striking the chip produces a cascade of high-energy phonons that rapidly propagate throughout the substrate. When the phonons collide with superconducting films of the qubits, they break Cooper pairs, producing quasiparticle excitations (QPs). When the QPs tunnel across the Josephson junction, they can exchange energy with the qubit degree of freedom. This results in an elevated qubit relaxation rate. Only when the QPs recombine or get trapped does the error burst end and the processor performance recover. This prompts the search for strategies to either accelerate the phonon decay [martinis_saving_2021, iaia_phonon_2022] or suppress QP tunneling.
@@
In this paper, we address this question by measuring QP bursts in transmons with different pairs of junction film thicknesses and, consequently, with different gap profiles. All of our transmons are made of aluminum, have 3D geometry [paik_observation_2011], and are deposited on sapphire substrates. First, we benchmark our gap engineering by measuring the gap difference, $\delta\Delta$, for our devices with different film thicknesses. The gap difference is obtained by monitoring the tunneling rate of resident QPs at different fridge temperatures. By varying the film thickness, we control the gap difference from 0 to 10 $\mathrm{GHz}$ (in frequency units). We then evaluate the effect of gap difference on the detection rate of burst events and measure qubit relaxation rates during bursts.
@@
We find that gap engineering reduces the burst detection rate by a factor of five. However, the reduction falls short of the four orders of magnitude suppression observed for the resident QPs. We believe that this discrepancy results from the elevation of the effective chip temperature during the burst to about 90 mK. The additional thermal energy allows the QPs to overcome the gap difference and tunnel across the junction, thus facilitating qubit relaxation. We directly confirm the persistence of elevated substrate temperature by comparing the qubit excitation and relaxation rates during the burst.
@@
Figure 1: (a) Parity-switching rate ($\Gamma$) as a function of the fridge temperature ($T$) for devices with different film thicknesses. The film thickness is used as a knob to control the gap difference $\delta\Delta$ at the qubit junction. Down (up)-pointing triangles represent rate $\Gamma_{0}$ ($\Gamma_{1}$) conditioned on the qubit residing in the ground (excited) state. When $\delta\Delta>hf_{q}$, the gap difference is efficient in obstructing the tunneling of resident QPs. Correspondingly, $\Gamma_{1}$ follows an Arrhenius law with the activation exponent $\delta\Delta-hf_{q}$, i.e., $\Gamma_{1}\propto\exp\left(-\left[\>\delta\Delta-hf_{q}\right]/\>k_{\mathrm{B}}T\right)$. This regime is realized in the big- and medium-$\delta\Delta$ device (left and middle panels). For the small-$\delta\Delta$ device (right panel), $\delta\Delta<hf_{q}$, and the gap difference is not effective at preventing the tunneling of resident QPs. The observed rate $\Gamma_{1}$ for the small-$\delta\Delta$ device is the highest among the measured devices. Comparing the data for $\Gamma_{1}$ and $\Gamma_{0}$ to theory (solid lines) allows us to extract $\delta\Delta$ and the resident QP density for all three devices. The resulting fit parameters are shown in Table 1. Notably, $\Gamma^{\mathrm{ph}}$ and $x_{\mathrm{QP}}^{\mathrm{ne}}$ for the three devices are of the same order of magnitude despite being fabricated individually. We attribute this consistency to the shared electromagnetic and radiation environment provided by the multiplexed package [noauthor_see_nodate]. (b) Exponential suppression of the resident QP tunneling rate, $\Gamma_{1}-\Gamma_{1}^{\mathrm{ph}}$, in state $|1\rangle$. Different points correspond to devices with different $\delta\Delta-\>hf_{q}$, different cooldowns, and temperatures [noauthor_see_nodate]. For convenience of comparison between different devices, the rate is normalized by the resident QP density, $\gamma^{\mathrm{QP}}_{1}\coloneqq(\Gamma_{1}-\Gamma_{1}^{\mathrm{ph}})/\>x_{\mathrm{QP}}^{\mathrm{ne}}$. When the gap difference surpasses the qubit energy, the normalized tunneling rate decreases exponentially with increasing $(\delta\Delta-hf_{q})/k_{\mathrm{B}}T$. The inset shows the full dependence of $\gamma^{\mathrm{QP}}_{1}$ on $(\delta\Delta-hf_{q})/k_{\mathrm{B}}T$. The normalized rate diverges when $\delta\Delta=\>hf_{q}$ but becomes temperature-insensitive for $\delta\Delta<\>hf_{q}$. The gray-shaded regions in the main plot and inset represent the theoretical prediction for the rate for a range of additional parameters such as $E_{J}$ or $E_{C}$.
@@
Our results should be contrasted with Ref. [mcewen_resisting_2024], where the gap difference – obtained by film thicknesses similar to our gap-engineered device – resulted in essentially no bursts being detected. The difference may arise from a combination of several factors. First, unlike ours, the chip in Ref. [mcewen_resisting_2024] has most of its area covered by the aluminum ground plane. Second, their substrate is silicon whereas ours is sapphire. Third, the chips are thermalized differently. Our chips are anchored to the package only by two isolated clamps each [noauthor_see_nodate], whereas Ref. [mcewen_resisting_2024] presumably used numerous wirebonds that anchor their chip. Finally, unlike in our measurement, Ref. [mcewen_resisting_2024] probed the qubit only sparsely, once every 100 $\mathrm{\mu s}$. Such a measurement may miss bursts recovering faster than 100 $\mathrm{\mu s}$.
@@
Experimental setup.— Our goal is to evaluate the effectiveness of gap engineering in suppressing decoherence caused by the QP bursts. The gap difference between the junction leads, with gaps of $\Delta$ and $\Delta+\>\delta\Delta$, prevents the tunneling of low-energy QPs from the low-gap lead to the high-gap one. The elevation of the superconducting gap in one of the leads, $\delta\Delta$, is controlled by changing the film thicknesses employed in the junction fabrication [chubov_dependence_1969, yamamoto_parity_2006, court_energy_2007, diamond_distinguishing_2022, marchegiani_quasiparticles_2022, connolly_coexistence_2024]. We fabricate three 3D aluminum transmons using different film thickness pairs: [17 nm, 83 nm], [23 nm, 123 nm], and [85 nm, 106 nm] (measured by atomic force microscopy). Based on prior measurements [chubov_dependence_1969, yamamoto_parity_2006, court_energy_2007, marchegiani_quasiparticles_2022], the corresponding $\delta\Delta$ are expected to be approximately 10 GHz (“big”-$\delta\Delta$), 5 GHz (“medium”-$\delta\Delta$), and 0.5 GHz (“small”-$\delta\Delta$), respectively, in frequency units. Table 1 summarizes relevant qubit parameters.
@@
The transmons are designed to be in the offset-charge-sensitive regime [riste_millisecond_2013, serniak_hot_2018, serniak_direct_2019], with the charge dispersion on the order of MHz. This enables the detection of charge-parity-switching events, which are experimental signatures of QP tunneling. By measuring the parity-switching rates, we can monitor how frequently QPs cross the junction in the presence of $\delta\Delta$. All three chips are mounted together in a multiplexed package, ensuring that electromagnetic environment is the same [noauthor_see_nodate].
@@
Table 1: Fit parameters and uncertainties corresponding to devices in Fig. 1(a). For the medium- and small-$\delta\Delta$ devices, the parity-switching rate is dominated by the QP tunneling rate, making it difficult to extract $\Gamma_{1}^{\mathrm{ph}}$. Therefore, for these devices, we assume $\Gamma_{0}^{\mathrm{ph}}=\Gamma_{1}^{\mathrm{ph}}$. In contrast, we can confidently fit $\Gamma_{1}^{\mathrm{ph}}$ for the big-$\delta\Delta$ device. This value can be found in [noauthor_see_nodate].
$f_{q}$ $\delta\Delta/h$ $\Delta/h$ $x_{\mathrm{QP}}^{\mathrm{ne}}$ $\Gamma_{0}^{\mathrm{ph}}$
(GHz) (GHz) (GHz) ($\times 10^{-10}$) ($\mathrm{s}^{-1}$)
[…]
@@
Measuring the gap difference.—As an initial step, we characterize the superconducting gap barriers, $\delta\Delta$, of our devices. To do this, we measure the parity-switching rates in the steady state, i.e., away from rare QP bursts. In this case, the parity-switching rates are dominated by the tunneling of resident non-equilibrium QPs populating superconducting films. The energy distribution of these QPs equilibrates with the base stage of the refrigerator [connolly_coexistence_2024]. The temperature dependence of the QP tunneling rate then follows the Arrhenius law, with the activation exponent controlled by $\delta\Delta$. Indeed, $\delta\Delta$ impedes the QPs across the junction, such that only few energetic QPs tunnel through the junction. Fitting the measured temperature dependence of the parity-switching rates to the Arrhenius law thus directly yields $\delta\Delta$, as we explain below. This measurement resembles that in Ref. [connolly_coexistence_2024]. Fig. 1(a) shows the measured parity-switching rates as a function of temperature for our three devices, and Fig. 1(b) shows a summary of additional devices and cooldowns.
@@
To fit the measured rates, we use a detailed theoretical model of QP tunneling in the presence of a gap difference at the junction [diamond_distinguishing_2022, marchegiani_quasiparticles_2022, connolly_coexistence_2024]. According to this theory, when the qubit is in the ground state, the parity-switching rate reads $\Gamma^{\mathrm{QP}}_{0}=\>\gamma^{\mathrm{QP}}_{0}\>x_{\mathrm{QP}}^{\mathrm{ne}}$. Here, $x_{\mathrm{QP}}^{\mathrm{ne}}$ is the temperature-independent density of resident QPs (in units of Cooper pair density), and $\gamma^{\mathrm{QP}}_{0}\propto E_{J}\exp\left(-\>\delta\Delta/\>k_{\mathrm{B}}T\right)$ is proportional to the tunneling rate and the fraction of QPs with energy exceeding $\delta\Delta$. Here, $k_{\mathrm{B}}$ is the Boltzmann constant, $T$ is the QP temperature, and $E_{J}$ is the Josephson energy of the junction. When the qubit is excited, and if $\delta\Delta>\>hf_{q}$, this expression modifies to $\Gamma^{\mathrm{QP}}_{1}=\>\gamma^{\mathrm{QP}}_{1}\>x_{\mathrm{QP}}^{\mathrm{ne}}$ with $\gamma^{\mathrm{QP}}_{1}\propto E_{J}\exp\left(-\left[\>\delta\Delta-\>hf_{q}\right]/\>k_{\mathrm{B}}T\right)$, where $h$ is the Planck constant and $f_{q}$ is the qubit frequency. The activation energy is reduced compared to that in $\gamma^{\mathrm{QP}}_{0}$ because the QP can absorb the excitation from the qubit, making it easier to overcome the gap difference. On contrary, if $\delta\Delta<\>hf_{q}$, then $\gamma^{\mathrm{QP}}_{1}$ becomes roughly temperature-independent. For each device, we simultaneously fit $\Gamma_{0}$ and $\Gamma_{1}$ to the expressions outlined above, using $\delta\Delta$ and $x_{\mathrm{QP}}^{\mathrm{ne}}$ as fitting parameters.
@@
There are two complications in this fitting protocol. First, at high temperatures, $T>100\>\mathrm{mK}$, the density of equilibrium QPs, $x_{\mathrm{QP}}^{\mathrm{th}}(T)=\sqrt{2\pi\>k_{\mathrm{B}}T/\Delta}\exp\left(-\Delta/\>k_{\mathrm{B}}T\right)$, becomes comparable to that of the resident non-equilibrium QPs, $x_{\mathrm{QP}}^{\mathrm{ne}}$. Therefore, at high temperatures, we account for the contribution of equilibrium QPs to the parity-switching rates. Second, there is a small background contribution $\Gamma^{\mathrm{ph}}$ to parity-switching rates coming from the absorption of stray pair-breaking photons at the junction [houzet_photon-assisted_2019, diamond_distinguishing_2022, connolly_coexistence_2024]. In our theoretical fit, we add such contributions as a temperature-independent offset. Notably, the contribution $\Gamma^{\mathrm{ph}}$ of the photons might differ for the two qubit states [houzet_photon-assisted_2019, diamond_distinguishing_2022]. We use notations $\Gamma_{0}^{\mathrm{ph}}$ and $\Gamma_{1}^{\mathrm{ph}}$ to differentiate this contribution between the respective qubit states.
[…]
@@
Burst measurement.—With the gap differences under control, we proceed to analyze the temporally correlated errors caused by QP bursts in the devices. To assess the effect of the gap difference on mitigating the high-energy impacts, we compare the bursts in the big- and small-$\delta\Delta$ devices. The results for the medium-$\delta\Delta$ device are discussed separately in [noauthor_see_nodate], due to the complications arising from its gap difference being comparable to the qubit energy.
@@
To detect the QP bursts, we attempt to stabilize the qubit in its excited state [mcewen_resolving_2022, harrington_synchronous_2025, dominicis_evaluating_2024, wu_mitigating_2025]. To this end, we repeatedly readout the qubit state every 5.7 $\mathrm{\mu s}$. If we find the qubit in its ground state $|0\rangle$, we apply a $\pi$-pulse to bring it to the excited state $|1\rangle$. The pulse sequence is illustrated in Fig. 2(b). During QP burst events, the qubit relaxation rate significantly increases. This allows us to identify the bursts by the excess number of observed $|1\rangle\rightarrow|0\rangle$ relaxation events [e.g. Fig. 2(d)]. In our measurement, we post-select events that do not cause large changes to the offset charge. However, we find that this post-selection is inconsequential for the burst event statistics [noauthor_see_nodate].
To collect sufficient statistics, we repeat the measurement for about 10 hours for each device. We then chop the measurement traces into $1\>\mathrm{ms}$ windows and tally the number of observed relaxation events in each window. Normalized histograms of the relaxation counts are shown in Fig. 2(c). For each device, the histogram contains a large peak (10 - 15 relaxation events), which is unrelated to the QP bursts. Although the peak of the distribution is described by the uncorrelated Poissonian relaxation events, there is an anomolous non-Poissonian tail containing an excess number of relaxation events. This tail stems from bursts containing correlated errors.
Its clear distinction from the Poissonian peak makes us confident that a burst has occurred within a $1\>\mathrm{ms}$ window when the number of relaxation events exceeded a certain threshold. In practice, we set the threshold at eight standard deviations away from the mean of both Poissonian distributions to make the false positive event rate negligible. Notably, some of the identified events are unrelated to bursts; they stem from unstable lossy defects in the environment [muller_interacting_2015, klimov_fluctuations_2018, burnett_decoherence_2019, muller_towards_2019, schlor_correlating_2019, lisenfeld_electric_2019]. These events can be effectively filtered out since their time dynamics is sufficiently different from that of the bursts [noauthor_see_nodate].